OpenStudy (anonymous):

Algbra 1 Question in reply.

OpenStudy (anonymous):

AC/BC is equal to ? A. ac/b B. c/c C. a/b D. bc/ac

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

?

OpenStudy (anonymous):

cancel the common factor from the numerator/denominator and what do you have left?

OpenStudy (anonymous):

as in \(\frac{5\times 11}{4\times 11}\)